OpenStudy (anonymous):

Find the zeros of the polynomial function. f(x) = x2 + 9x + 20 A. 2, 10 B. –4, –5 C. –2, –10 D. 4, 5

OpenStudy (anonymous):

@grace111

OpenStudy (the_fizicx99):

Find the zeros :3

OpenStudy (the_fizicx99):

0 = x^2 + 9x + 20 4 * 5 = 20 4 + 5 = 9 0 = (x+4)(x+5) x + 4 = 0 --> x = -4 x + 5 = 0 --> x = -5

OpenStudy (anonymous):

So, B?

OpenStudy (the_fizicx99):

Yes, when you plug in those values you'll get 0

OpenStudy (anonymous):

Thanks :D

OpenStudy (the_fizicx99):

Yw :D